OpenStudy (anonymous):

A theater company has raised $810.75 by selling 15 floor seat tickets. Each ticket costs the same. Part A: Write an equation with a variable that can be solved to correctly find the price of each ticket. Explain how you created this equation. (5 points) Part B: Solve your equation in Part A to find the price of each floor seat ticket. How do you know your solution is correct? (5 points)

OpenStudy (anonymous):

Part A would be $810.75/15=n (I think...) Part B would be $810.75/15= $54.05